c. Why should interest rates be generally positive? (3 marks) d. Under what circumstances will the NPV rule give the same decision as the IRR rule? (5 marks) Luther Industries just paid dividends of $4.50 and has a cost of equity capital of 10%. Luther Industries dividends are expected to grow at a constant rate of 4% pa indefinitely. What is the price of Luther Industries shares? e. (5 marks)
Answer #1

c

Interest rates should be able to attract people to invest more in the economy. If they are positive people will invest in the economy and this will help the economy to grow. If they are negative, people will lose money and will not invest.

d

NPV rule will be the same as IRR rule when the projects are independent and cash flows are positive that is they are conventional cash flows. Also the life of projects should be the same with a single initial investment in the project.

e

Share price= D1/ (k-g) = D0*(1+g)/ (k-g)

= 4.5*(1+0.04)/ (0.1-0.04)

=$78